paul is 4 years younger than his brother. in four years, paul will be three-fourths as old as his brother. how old are they now?
Answer by JBarnum(2146)

Paul Brother
P=B-4
P+4=(3/4)(B+4)
B-4+4=(3/4)B+3
(1/4)B=3
B=12
P=8